One leg of a right triangle is 5 cm shorter than the other leg. The length of the hypotenuse is 25 cm. Find the length of the shorter leg.
Answer by mananth(16946)

let one leg be x cm
the other leg will be x-5 cm
hyptenuse = 25 cm
Using pytagorean theorem
x^2 +(x-5)^2= 25^2
X^2 +x^2-10x+25= 625
2x^2 -10x -600=0
x^2-5x-300=0
x^2-20x+15x-300=0
x(x-20)+15(x-20)=0
(x-20(x+15)=0
x=20 which is positive
one leg is 20cm
the other leg = 15 cm